Given Q=400-8P + 0.05Y. Where P=15 and Y=200. Fina the price elasticity of demand and income elasticity of demand

1
Expert's answer
2021-11-18T11:41:44-0500
"Q=400-8\times15+0.05\times200=290"

"E_p=-8\times\frac{15}{290}=-0.41"

"E_Y=0.05\times\frac{200}{290}=0.03"